# SentenceVAE: Faster, Longer and More Accurate Inference with Next-sentence Prediction for Large Language Models

### v1

@@ Abstract
Contemporary large language models (LLMs) predominantly utilize a next-token prediction method for inference, which significantly impedes their processing speed. In this paper, we introduce a novel inference methodology termed next-sentence prediction, aimed at enhancing the inference efficiency of LLMs. We present SentenceVAE, a tiny model consisting of an encoder and a decoder. The encoder effectively condenses the information within a sentence into a singular token, while the decoder reconstructs this compressed data back into its original sentential form. By integrating SentenceVAE into the input and output layers of LLMs, we develop Sentence-level LLMs (SLLMs) that employ a sentence-by-sentence inference approach, markedly accelerating inference speeds. SentenceVAE also maintains the integrity of the original semantic content by segmenting the text into sentences, thereby preserving accuracy while boosting inference speeds. Compared to traditional LLMs, SLLMs process fewer tokens over equivalent context lengths, significantly reducing memory demands for Self-Attention computations and facilitating the handling of longer contexts. Our experimental findings reveal that this method can increase inference speeds by 204$\sim$365%, reduce perplexity (PPL) to 46$\sim$75of its original metric, and decrease memory overhead by 86$\sim$91for the same context length. The advantages of this approach are further amplified with increases in model parameters.
@@ Introduction
Large language models (LLMs) have demonstrated remarkable proficiency in understanding human intent through the acquisition of knowledge from vast and diverse datasets. These models provide rich and accurate information and responses, making them indispensable in applications such as machine translation (Zhang, Haddow, and Birch 2023; Wang et al. 2023), chatbots (Islam and Moushi 2024; Wang et al. 2024), and question-answering systems (Singhal et al. 2023; Lazaridou et al. 2022). The predominant approach to inference in LLMs involves next-token prediction, where the model generates tokens sequentially. While this token-by-token generation method has achieved significant performance, it is inherently limited in that it produces only one token per inference step. This limitation results in substantial computational overhead and increases the overall inference time, thereby constraining the scalability and efficiency of LLMs. As a result, there is a pressing need to explore alternative methods that can mitigate these limitations while maintaining or improving upon the accuracy and responsiveness of LLMs.
@@
To enhance the inference efficiency of Large Language Models (LLMs) and reduce computational costs, Gloeckle et al. (Gloeckle et al. 2024) proposed training LLMs to predict multiple tokens simultaneously, aiming to accelerate the inference process. However, the number of tokens predicted at once is fixed. For instance, when this number is set to five, regardless of the input text, it is forcibly segmented into units of five tokens. While this multi-token prediction method improves the inference speed of LLMs, it may disrupt the original semantics of the text by aggregating unrelated tokens together, potentially compromising inference accuracy.
To enhance inference speed while maintaining accuracy, we explore methods for adaptively selecting the optimal number of predicted tokens based on the semantic content of the text. Consequently, we introduce SentenceVAE, which comprises a sentence encoder and a sentence decoder. The schematic of SentenceVAE is illustrated in Fig. 1. The sentence encoder is designed to condense the information of an entire sentence into a single token, while the sentence decoder reconstructs this compressed token back into a sentence. By integrating the encoder and decoder of SentenceVAE into the input and output layers of Large Language Models (LLMs), these models can perform next-sentence prediction, thereby accelerating the reasoning process. Moreover, since SentenceVAE segments text at the sentence level, the semantic integrity is better preserved, ensuring the accuracy of the inferrence.
Specifically, prior to inputting text into LLMs, SentenceVAE segments the text into sentences and utilizes the sentence encoder to compress the information contained within each sentence into a single token. These compressed tokens are then fed as input to the LLMs. The LLMs predict their output based on these compressed tokens. Finally, the sentence decoder decodes the predicted token into the final output of the LLMs.
@@
  • We propose SentenceVAE, a model capable of compressing the information content of a sentence into a single token, and conversely, expanding such tokens to reconstruct sentences with high informational fidelity.
@@
Figure 2: (a) The schematic form of traditional LLM. (b) The schematic form of Setence-level LLM (SLLM), which embedded with SentenceVAE. It can be clearly seen that, unlike the next-token inference method employed by traditional large language models (LLMs), the proposed method adopts a next-sentence prediction approach, which significantly reduces the number of inference iterations and the overall inference cost of LLMs.
@@
Large language models: In recent years, large language models (LLMs) have achieved remarkable performance in the field of natural language processing (NLP). However, these models typically possess a vast number of parameters from billions to hundreds of billions, exemplified by models such as TeleChat (1B-52B) from TeleAI (Wang et al. 2024; Li et al. 2024b), Llama (7B-405B) from Meta (Touvron et al. 2023a, b), InternLM (1.8B-20B) from Shanghai AI Lab (Cai et al. 2024), etc. The sheer magnitude of these parameters renders the inference process computationally intensive and time-consuming. Moreover, most LLMs employ a one-token prediction method, where the model predicts only one token at a time. This token-by-token approach further exacerbates the inference time of LLMs. As a result, accelerating the inference speed of LLMs while maintaining their accuracy presents a significant challenge for researchers in the field.
@@
Multi token prediction: The transformer model enables the parallel processing of sequences through its self-attention mechanism, providing a technical foundation for the implementation of multi-token prediction (Vaswani et al. 2017). The BERT model utilizes bidirectional encoder representations to capture contextual information, further solidifying the basis for the implementation of multi-token prediction (Devlin et al. 2018). The T5 model incorporates multi-token prediction during the training process for the first time, significantly enhancing the coherence and quality of the generated text (Raffel et al. 2020). The Better & Faster LLMs apply multi-token prediction to LLMs, allowing them to directly predict multiple tokens in a single inference step, thereby reducing the number of inference iterations and associated computational costs (Gloeckle et al. 2024). However, the number of predicted tokens is fixed, which may inadvertently combine unrelated tokens, potentially disrupting the original semantics of the sentence and consequently decreasing inference accuracy.
@@
Encoder-decoder model: The concept of an encoder is introduced by Yann LeCun in his doctoral dissertation, wherein he posited that an encoder could transform input data into shorter representations without losing the core information, thereby providing a foundational approach to compressing high-dimensional data. The Variational Autoencoder (VAE) model introduces the use of hidden variables between the encoder and decoder to learn the data distribution, laying the groundwork for deep-learning-based encoders and decoders (Kingma and Welling 2013). Sutskever et al. introduces the encoder-decoder framework into the field of deep learning, utilizing recurrent neural networks (RNNs) as both the encoder and decoder for sequence-to-sequence (Seq2Seq) learning, demonstrating exceptional performance in machine translation tasks (Sutskever, Vinyals, and Le 2014). The SegNet applies the encoder-decoder model to the field of image segmentation, significantly enhancing segmentation performance (Badrinarayanan, Kendall, and Cipolla 2017). Subsequently, the encoder-decoder model has achieved great success in various domains including machine translation (Cho et al. 2014; Makin, Moses, and Chang 2020), text-to-speech conversion (Li et al. 2020), computational imaging (Chen et al. 2023; Li et al. 2024a), and beyond, fully demonstrating the capability of the encoder-decoder architecture to efficiently compress and restore data.
@@
By leveraging the advantages of multi-token prediction and the encoder-decoder model, we propose SentenceVAE. SentenceVAE is capable of compressing the information contained within a sentence into a single token, thereby enabling next-sentence prediction and significantly accelerating the inference speed of LLMs. Additionally, This approach divides the text into sentences without compromising the original semantics, thereby accelerating the inference speed of LLMs while ensuring inference accuracy.
@@
In this section, we introduce the SentenceVAE framework. Embedding this framework into a LLM enables the model to implement a sentence-by-sentence prediction method, significantly accelerating its inference speed. The inference process of the proposed method compared to traditional LLMs is illustrated in Fig. 2. It is evident that for the same text, an LLM embedded with SentenceVAE requires only three inference steps to obtain the inference result, whereas traditional LLMs necessitate twenty-free inference steps. Furthermore, the proposed method divides the text into sentences without compromising the original semantics, thereby enhancing the inference speed of LLMs while maintaining inference accuracy.
@@
Our SentenceVAE primarily consists of a Sentence Encoder and a Sentence Decoder. The Sentence Encoder encodes multiple word-level tokens from a sentence into a single sentence-level token, while the Sentence Decoder reconstructs this sentence-level token back into the original sequence of word-level tokens. To ensure that the multiple tokens input to the Sentence Encoder come from individual sentences, we propose a specialized Sentence Segmentation Mechanism. Additionally, we introduce a Feature Fusion Mechanism, which enable the encoder to encode variable-length token sequences into a single sentence-level token.
@@
Sentence Encoder: Taking $s=s_{t}(1\leq t\leq n)$ as an example, after tokenization, we obtained a sequence of $L$ token ids $\boldsymbol{D}=[d_{1},d_{2},d_{3},...,d_{L}]$ representing a sentence. This sequence was then passed through an embedding layer, resulting in word-level embedding vectors of $\boldsymbol{E}$, as shown in Eq.1.
@@
These embeddings were subsequently input into self-attention based encoder blocks, yielding hidden features $\boldsymbol{H}$, as shown in Eq. 2.
@@
To derive the sentence embedding vector, we fused these $L$ features into a single vector $\boldsymbol{\Omega}_{t}\in\mathbb{R}^{\text{hidden\_size}}$. This sentence embedding vector is then fed into the decoder-only LLM, which generates a new sentence embedding vector $\boldsymbol{\Omega}_{t+1}$ at each time step.
@@
In this section, we introduce a unified grafting method to graft SentienceVAE onto the beginning and end of any LLM. This approach transforms the LLMs into Sentence-level Large Language Models (SLLMs), enabling it to operate effectively in the sentence-level embedding space.
@@
Termination judgment layer: In traditional LLMs, the final fully connected layer $\textbf{FC}_{\text{llm}}$ typically converts an $\boldsymbol{H}_{\text{llm}}\in\mathbb{R}^{t\times\text{hidden\_size}}$ to a probability vector (logits) $\boldsymbol{P}_{\text{llm}}\in\mathbb{R}^{t\times V}$, where $V$ represents the size of the vocabulary. This allows the model to determine the next token output using sampling algorithms, e.g., Greedy Search, Beam Search (Graves 2012), Random Sampling, Temperature Sampling (Hinton, Vinyals, and Dean 2015), Top-K Sampling (Fan, Lewis, and Dauphin 2018), and Nucleus Sampling (Holtzman et al. 2019). If the current token output is a special token like eos (end-of-sequence), the iteration terminates.
However, in SLLMs, the output of $\textbf{DB}_{\text{llm}}$ is a sentence-level hidden state vector $\boldsymbol{H}_{\text{sllm}}$. Therefore, traditional token generation methods are not applicable. Instead, we use a new fully connected layer $\textbf{FC}_{\text{sllm}}$ called termination judgment layer that converts the $\boldsymbol{H}_{\text{sllm}}$ into a 2-dimensional boolean vector $\boldsymbol{B}_{\text{stop\_flag}}$, as shown in Eq. 8. This vector helps determine whether the current sentence-level hidden state vector signifies the end of a sentence (stop flag) or needs further decoding by the sentence decoder. If the vector indicates an end flag, the iteration terminates. Otherwise, the sentence decoder processes the embedding to generate corresponding tokens, as shown in Eq. 9.
@@
Inferencing sentence by sentence: In the SLLM architecture, we use a small model (SentenceVAE) to encode multiple tokens within a sentence into a single sentence embedding vector. This enables the LLM to conduct inference on a sentence-by-sentence basis rather than the traditional token-by-token approach. Experimental results indicate that this approach can significantly reduce the computational cost required for model inference and enhance inference speed.
@@
ID Input Sentence Toks Output Sentence

### v3

# SentenceVAE: Enable Next-sentence Prediction for Large Language Models with Faster Speed, Higher Accuracy and Longer Context

@@ Abstract
Current large language models (LLMs) primarily utilize next-token prediction method for inference, which significantly impedes their processing speed. In this paper, we introduce a novel inference methodology termed next-sentence prediction, aimed at enhancing the inference efficiency of LLMs. We present Sentence Variational Autoencoder (SentenceVAE), a tiny model consisting of a Sentence Encoder and a Sentence Decoder. The Sentence Encoder can effectively condense the information within a sentence into a singular token, while the Sentence Decoder can reconstruct this compressed token back into sentence. By integrating SentenceVAE into the input and output layers of LLMs, we develop Sentence-level LLMs (SLLMs) that employ a sentence-by-sentence inference method. In addition, the SentenceVAE module of SLLMs can maintain the integrity of the original semantic content by segmenting the context into sentences, thereby improving accuracy while boosting inference speed. Moreover, compared to previous LLMs, SLLMs process fewer tokens over equivalent context length, significantly reducing memory demands for self-attention computation and facilitating the handling of longer context. Extensive experiments on Wanjuan dataset have reveal that the proposed method can accelerate inference speed by 204$\sim$365%, reduce perplexity (PPL) to 46$\sim$75of its original metric, and decrease memory overhead by 86$\sim$91for the equivalent context length, compared to the token-by-token method.
@@ Introduction
Large language models (LLMs) have demonstrated remarkable proficiency in understanding human intent through the acquisition of knowledge from vast and diverse datasets. These models provide rich and accurate responses, making them essential for applications such as machine translation (Zhang, Haddow, and Birch 2023; Wang et al. 2023), chatbots (Islam and Moushi 2024; Wang et al. 2024), and question-answering systems (Singhal et al. 2023; Lazaridou et al. 2022). The predominant method to inference in LLMs involves next-token prediction, where the model generates tokens sequentially. While this token-by-token generation method has achieved significant performance, it is inherently limited in producing only one token per inference step. This constraint introduces considerable computational overhead, prolonging the inference process and impeding the scalability and operational efficiency of LLMs. Consequently, there is a critical need to explore alternative methods that can improve the inference speed while preserving or even enhancing the accuracy and responsiveness of LLMs.
@@
To enhance the inference speed of LLMs, Gloeckle et al. (Gloeckle et al. 2024) proposed training LLMs to predict multiple tokens simultaneously, aiming to accelerate the inference process. However, the number of tokens predicted simultaneously is fixed, resulting in a rigid and inflexible partitioning of the input context. For example, when set to five, the input is invariably segmented into non-overlapping sequences of five tokens, regardless of the context’s underlying structure, potentially compromising inference accuracy.
We revisited the mechanisms underlying human interaction. Typically, when individuals prepare to speak, they have already formulated the semantic content of an entire sentence in their minds. However, due to the limitations of the vocal cords, which are relatively inefficient organs for sound production, speakers are compelled to articulate sentences word-by-word, much like the token-by-token generation method of LLMs. Similarly, listeners tend to comprehend sentences holistically, rather than word-by-word. This observation suggests that sentences, rather than individual words, constitute the fundamental units of comprehension in natural language. Motivated by this insight, we developed the Sentence Variational Autoencoder (SentenceVAE), as depicted in Fig. 1. This model comprises a Sentence Encoder and a Sentence Decoder. The encoder compresses the multiple tokens of a sentence into a single sentence-level token, mirroring the processing capabilities of the human auditory system, while the decoder reconstructs the sentence’s semantic content from this token, analogous to the function of the vocal cords. By integrating the encoder and decoder components of SentenceVAE seamlessly into the input and output layers of LLMs, these enhanced models are capable of operating within a more efficient sentence-level embedding space, performing next-sentence prediction.
@@
  • We propose SentenceVAE, a model capable of compressing the information content of a sentence into a single token and expanding such tokens to reconstruct sentences with high informational fidelity.
@@
Figure 2: (a) The schematic form of published LLMs. (b) The schematic form of SLLMs, which embedded with SentenceVAEs. It can be clearly seen that, unlike the next-token inference method employed by published LLMs, the proposed method adopts a next-sentence prediction method, which significantly reduces the number of inference iterations and the overall inference cost.
@@
Large language models: recently, LLMs have achieved remarkable performance in the field of natural language processing (NLP). However, these models typically possess a vast number of parameters from billions to hundreds of billions, exemplified by models such as TeleChat (1B-52B) from TeleAI (Wang et al. 2024; Li et al. 2024b), Llama (7B-405B) from Meta (Touvron et al. 2023a, b), InternLM (1.8B-20B) from Shanghai AI Lab (Cai et al. 2024), etc. The sheer magnitude of these parameters renders the inference process computationally intensive and time-consuming. Moreover, most LLMs employ a one-token prediction method, where the model predicts only one token at a time. This token-by-token method further exacerbates the inference time of LLMs. As a result, accelerating the inference speed of LLMs while preserving or enhancing their accuracy presents a significant challenge for researchers.
@@
Multi token prediction: the transformer model enables the parallel processing of sequences through its self-attention mechanism, providing a technical foundation for the implementation of multi-token prediction (Vaswani et al. 2017). The BERT model utilizes bidirectional encoder representations to capture contextual information, further solidifying the basis for the implementation of multi-token prediction (Devlin et al. 2018). The T5 model incorporates multi-token prediction during the training process for the first time, significantly enhancing the coherence and quality of the generated text (Raffel et al. 2020). The Better & Faster LLMs apply multi-token prediction to LLMs, allowing them to directly predict multiple tokens in a single inference step, thereby reducing the number of inference iterations and associated computational costs (Gloeckle et al. 2024). However, the number of predicted tokens is fixed, which may inadvertently combine unrelated tokens. This can result in the merging of unrelated tokens, disrupting the original sentence structure and consequently reducing inference accuracy.
@@
Encoder-decoder model: the concept of an encoder is introduced by Yann LeCun in his doctoral dissertation (Le Cun 1987), wherein he posited that an encoder could transform input data into shorter representations without losing the core information, thereby providing a foundational approach to compressing high-dimensional data. The Variational Autoencoder (VAE) model introduces the use of hidden variables between the encoder and decoder to learn the data distribution, laying the groundwork for deep-learning-based encoders and decoders (Kingma and Welling 2013). Sutskever et al. introduces the encoder-decoder framework into the field of deep learning, utilizing recurrent neural networks (RNNs) as both the encoder and decoder for sequence-to-sequence (Seq2Seq) learning, demonstrating exceptional performance in machine translation tasks (Sutskever, Vinyals, and Le 2014). The SegNet applies the encoder-decoder model to the field of image segmentation, significantly enhancing segmentation performance (Badrinarayanan, Kendall, and Cipolla 2017). Subsequently, the encoder-decoder model has achieved great success in various domains including machine translation (Cho et al. 2014; Makin, Moses, and Chang 2020), text-to-speech conversion (Li et al. 2020), computational imaging (Chen et al. 2023; Li et al. 2024a), and beyond, fully demonstrating the capability of the encoder-decoder architecture to efficiently compress and restore data.
@@
By leveraging the advantages of multi-token prediction and the encoder-decoder model, we propose SentenceVAE. SentenceVAE is capable of compressing the information contained within a sentence into a single token, enabling next-sentence prediction and significantly accelerating the inference speed of LLMs. Additionally, our method divides the context into sentences without compromising the original semantics, accelerating the inference speed of LLMs while ensuring inference accuracy.
@@
In this section, we introduce the SentenceVAE framework. Embedding this framework into a LLM enables the model to implement a sentence-by-sentence prediction method, significantly accelerating its inference speed. The inference process of the proposed method compared to published LLMs is illustrated in Fig. 2. It is evident that for the same context, an LLM embedded with SentenceVAE requires only three inference steps to obtain the inference result, whereas previous LLMs necessitate twenty-three inference steps. Furthermore, the proposed method divides the context into sentences without compromising the original semantics, thereby enhancing the inference speed of LLMs while preserving or even enhancing accuracy.
@@
Our SentenceVAE primarily consists of a Sentence Encoder and a Sentence Decoder. The ecoder encodes multiple word-level tokens from a sentence into a single sentence-level token, while the decoder reconstructs this sentence-level token back into the original sequence of word-level tokens. To ensure that the multiple tokens input to the encoder come from individual sentences, we propose a specialized sentence segmentation mechanism. Additionally, we introduce a feature fusion mechanism, which enable the encoder to encode variable-length token sequences into a single sentence-level token.
@@
Sentence Encoder: taking $s=s_{t}(1\leq t\leq n)$ as an example, after tokenization, we obtaine a sequence of $L$ token ids $\boldsymbol{D}=[d_{1},d_{2},d_{3},...,d_{L}]$ representing a sentence. This sequence is then passed through an embedding layer, resulting in word-level embedding vectors of $\boldsymbol{E}$, as shown in Eq.1.
@@
These embeddings are subsequently input into self-attention based encoder blocks, yielding hidden features $\boldsymbol{H}$, as shown in Eq. 2.
@@
To derive the sentence embedding vector, we fuse these $L$ features into a single vector $\boldsymbol{\Omega}_{t}\in\mathbb{R}^{\text{hidden\_size}}$. This sentence embedding vector is then fed into the decoder-only LLM, which generates a new sentence embedding vector $\boldsymbol{\Omega}_{t+1}$ at each time step.
@@
In this section, we introduce a unified grafting method to graft SentienceVAE onto the beginning and end of any LLM. This method transforms the LLMs into SLLMs, enabling it to operate effectively in the sentence-level embedding space.
@@
Termination judgment layer: in previous LLMs, the final fully connected layer $\textbf{FC}_{\text{llm}}$ typically converts an $\boldsymbol{H}_{\text{llm}}\in\mathbb{R}^{t\times\text{hidden\_size}}$ to a probability vector (logits) $\boldsymbol{P}_{\text{llm}}\in\mathbb{R}^{t\times V}$, where $V$ represents the size of the vocabulary. This allows the model to determine the next token output using sampling algorithms, e.g., Greedy Search, Beam Search (Graves 2012), Random Sampling, Temperature Sampling (Hinton, Vinyals, and Dean 2015), Top-K Sampling (Fan, Lewis, and Dauphin 2018), and Nucleus Sampling (Holtzman et al. 2019). If the current token output is a special token like eos (end-of-sequence), the iteration terminates. Notably, in SLLMs, the output of $\textbf{DB}_{\text{llm}}$ is a sentence-level hidden state vector $\boldsymbol{H}_{\text{sllm}}$. Therefore, traditional token generation methods are not applicable. Instead, we use a new fully connected layer $\textbf{FC}_{\text{sllm}}$ called termination judgment layer that converts the $\boldsymbol{H}_{\text{sllm}}$ into a 2-dimensional boolean vector $\boldsymbol{B}_{\text{stop\_flag}}$, as shown in Eq. 8. This vector helps determine whether the current sentence-level hidden state vector signifies the end of a sentence (stop flag) or needs further decoding by the sentence decoder. If the vector indicates an end flag, the iteration terminates. Otherwise, the sentence decoder processes the embedding to generate corresponding tokens, as shown in Eq. 9.
@@
Inferencing sentence by sentence: in the SLLM architecture, we use a tiny SentenceVAE model to encode multiple tokens within a sentence into a single sentence embedding vector. This enables the LLM to conduct inference on a sentence-by-sentence basis rather than the traditional token-by-token method.
Experiment
To validate our method, we initially trained individual SentienceVAEs using self-supervised methods, demonstrating the capability of compressing multiple tokens from a sentence into a single vector via an encoder and reconstructing the original sequence through a decoder. Subsequently, we integrated these encoders and decoders at the endpoints of open-source LLMs, thereby enhancing LLMs into SLLMs that operate within a sentence embedding space. This modification not only improves perplexity (PPL) and enhances inference speed, but also reduces memory consumption. Observations of the loss curve indicate that SLLMs continue to adhere to the Scaling Law (Kaplan et al. 2020).
Experimental Setting
Our experiments utilized either a single 4-card RTX 4090 (24G) or a 4-card A100 (40G) (for SLLM-1.3B), employing data parallel distribution for training.
Base Models: we employed the 125M, 350M, and 1.3B models from the OPT series (Zhang et al. 2022) as our base LLMs, exploring the extension to larger models via the Scaling Law.
Dataset: the training dataset comprised the English subset (EN/WebText) of the Wanjuan-1.0 dataset (He et al. 2023), with SentenceVAE samples including approximately 153.6M sentences (1.7B tokens) and SLLM samples encompassing about 6.4M paragraphs (5.6B tokens). A validation set of 1,000 random sentences or paragraphs ensured no overlap with the training data.
Hyperparameters: for SentenceVAEs, we set the maximum tokens per sentence at 64, with a batch size of 128/card, a base learning rate of 1e-7/batch/card, and up to 300K iterations. SLLMs training mirrored these settings, with a maximum of 64 sentences per paragraph, a batch size of 1/card, a base learning rate of 1e-6/batch/card, and up to 1.6M iterations.
All experiments utilized the AdamW optimizer (Loshchilov and Hutter 2017) with AMP (Micikevicius et al. 2017) enabled, a weight decay coefficient of 0.01, and a maximum gradient L2 norm of 1 for clipping. The initial 5K iterations followed a linear learning rate schedule, transitioning to a cosine annealing schedule for subsequent iterations, alongside an Exponential Moving Average (EMA) strategy.
Metrics: PPL was employed as the evaluation metric, calculated for output logits $\boldsymbol{P}$ and ground truth $\boldsymbol{D}_{\text{groud\_truth}}$ as per the defined formula Eq. 10.
A series of word-level tokens can be represented by a sentence-level token with strong robustness
We trained SentienceVAEs using a self-supervised approach to assess if sentence embeddings could effectively represent and reconstruct word-level tokens. Models were trained with hidden sizes corresponding to the dimensions of the OPT-125M, OPT-350M, and OPT-1.3B, with varying block layers. Performance was evaluated using cross-entropy loss and PPL on the validation set, confirming our method. The experimental results are presented in Table 1.
@@
ID Input Sentence Tokens Output Sentence